35.) Show a complete solution to each problem
Find the exact length of the side of a square whose diagonal is 3 feet.

Answer by midwood_trail(310)

Find the exact length of the side of a square whose diagonal is 3 feet.
We are given the diagonal to be 3 feet but no side of the square is given.
Let x = side of the square.
Let 3 = the hypotenuse
We use the Pythagorean Theorem to find x.
x^2 + x^2 = 3^2
2x^2 = 9
Divide both sides by 2.
x^2 = 9/2
Take the square root of both sides.
x = sqrt{9/2}
x = sqrt{9}/sqrt{2}
x = 3/sqrt{2}
We rationalize the denominator.
Final answer:
x = 3(sqrt{2})/2